OpenStudy (anonymous):

what are the solutions to z^2 -6z-27=0

OpenStudy (anonymous):

-9, -3

OpenStudy (anonymous):

What have you done so far with the question @minimoo

OpenStudy (anonymous):

factored it out to be (z-9)(z+3)

OpenStudy (anonymous):

Yep that's good. Now you have this: \[(z-9)(z+3)=0\] What values of z make the equation zero?

OpenStudy (anonymous):

it would be z-9=0 and z+3=0 and the answers would be 9 and -3

OpenStudy (anonymous):

Well Done! That's correct.

OpenStudy (anonymous):

:) im getting the hang of it!

OpenStudy (anonymous):

Yep, you don't need to fret about this too much because if you try attempting it, you would be likely to get it right! ;)

OpenStudy (anonymous):

thanks again :)